A train travelling 72 km/h undergoes a uniform retardation (deceleration)of 2ms -2 when the breaks are applied. Find the time taken to come to rest and the distance travelled from the place where the breaks were applied


1
Expert's answer
2021-03-02T18:03:07-0500

(a)

"t=\dfrac{v-v_0}{a},""t=\dfrac{0-20\ \dfrac{m}{s}}{-2.0\ \dfrac{m}{s^2}}=10\ s."

(b)

"d=\dfrac{1}{2}(v+v_0)t,""d=\dfrac{1}{2}\cdot(0+20\ \dfrac{m}{s})\cdot10\ s=100\ m."